QUESTION 6

Answer the below questions :

A- PW for machine C =

B- Based on the PW value you got in the previous questions, which machine we should select? type you explanation below 
Answer #1

SOLUTION :


r = MARR = 10% = 0.1

=> (1 + r) = 1.1


PV of costs of Machine A :


= 15000 + 12340/ 1.1 + 12340/1.1^2 + 12340/1.1^3 - 4000/1.1^3 

= 42682.49 ($) 


PV of costs of Machine B :


= 30000 + 6000 (1.1^6 - 1)/(0.1*1.1^6) - 5000/1.1^6

= 53309.19 ($) 


PV of costs of Machine C :


= 10423 + 4000/1.1 + 4000/1.1^2 - 1000/1.1^2

= 16538.70 ($)


Therefore, as per PV of costs , Machine C has lowest cost. So Machine C should be selected.

(ANSWER)

answered by: Tulsiram Garg